Circular No. 6275

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995bc IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 18.5) at α =

9^{h}51^{m}43^{s}.50, δ = +40^{o}18′50′′.6 (equinox 2000.0), which is 5′′.0 east and 14′′.2 north of the galaxy’s center (positions measured by Mueller and reductions performed by B. Skiff and L. Wasserman, Lowell Observatory). The IIIa-J discovery plate was taken on Dec. 1 UT at the 1.2-m Oschin Schmidt Telescope by J. D. Mendenhall and Mueller in the course of the second Palomar Sky Survey. No object appears at the supernova position on original Sky Survey prints or on IIIa-F second Survey transparencies.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that CCD spectra (range 420–700 nm, resolution 0.7 nm) obtained on Dec. 17 UT with the 3-m Shane reflector at Lick Observatory show that SN 1995bc is a type-II supernova. There are prominent, broad hydrogen Balmer lines having P-Cyg profiles. The redshift of the parent galaxy, measured from narrow Hα and [N II] lines, is ∼ 0.048.

NEW X-RAY TRANSIENT REPEATER K. Hurley, on behalf of the Ulysses Gamma-Ray Burst Team; C.

Kouveliotou, A. Harmon, G. Fishman, and M. Briggs, on behalf of the BATSE/GRO team; J. van Paradijs, University of Amsterdam; and J. Kommers and W. Lewin, Massachusetts Institute of Technology, report: “A preliminary network synthesis position (cf. Hurley 1994, Ap.J. 431, L31) for the new repeating x-ray transient source (IAUC 6272) has been obtained using Ulysses and BATSE (specifically, 59 BATSE events from Dec. 2). The position is an annulus whose half-width is 0^{o}.21 centered at α = 2^{h}44^{m}37^{s}, δ = −55^{o}31′.0 (equinox 2000.0), with central radius 87^{o}.500. Using the BATSE occultation technique, we excluded a large area of the sky for the source origin. The intersection of the network synthesis annulus with the remaining (allowed) area of detection may be approximated by a box with the following coordinates: α = 17^{h}50^{m}, δ = −28^{o}.1; 17^{h}49^{m}, −28^{o}.5; 17^{h}23^{m}, −30^{o}.3; 17^{h}19^{m}, −31^{o}.0.”

NOVA CASSIOPEIAE 1995 Visual magnitude estimates: Nov. 25.79 UT, 8.4 (A. Pereira, Cabo

da Roca, Portugal); Dec. 3.99, 8.6 (P. Schmeer, Bischmisheim, Germany); 9.75, 8.0 (Schmeer); 14.75, 7.4 (B. H. Granslo, Fjellhamar, Norway); 15.74, 7.4 (A. Diepvens, Balen, Belgium); 16.74, 7.2 (E. Broens, Mol, Belgium).

1995 December 17 Daniel W. E. Green